# Fixing That Free Lunch: When, Where, and Why Synthetic Data Fails in Model-Based Policy Optimization

### v2

@@ 1 Introduction
The “no free lunch” principle highlights that no optimization method achieves universal optimality across every problem instance. In reinforcement learning (RL), this underscores that an algorithm’s effectiveness is shaped by the specific environment and its characteristics. For Dyna style model-based RL (MBRL) algorithms, wherein synthetic model-generated data is intended to improve sample efficiency and enable rapid policy improvement, recent evidence suggests that these benefits are not guaranteed. In Stealing That Free Lunch (STFL) (Barkley & Fridovich-Keil, 2025), Model Based Policy Optimization (MBPO) (Janner et al., 2019), a widely used and well cited Dyna style algorithm, was shown to match or exceed its model free counterpart, Soft Actor Critic (SAC) (Haarnoja et al., 2019), in OpenAI Gym (Brockman et al., 2016). However, in many DeepMind Control Suite (DMC) (Tassa et al., 2020) tasks, MBPO then failed to improve beyond a randomly initialized policy when trained from scratch, despite both benchmarks sharing the same MuJoCo physics engine (Todorov et al., 2012) and featuring similar high level task reward structures.
While STFL established that these failures exist, it did not explain why MBPO collapses in specific settings, other than noting the central role of synthetic data in driving failures. In this work, we close that gap. We focus on challenging DMC tasks where MBPO failed completely in STFL, often performing no better than a random initial policy throughout training (Barkley & Fridovich-Keil, 2025). Through targeted experiments, we identify two coupled mechanisms driving these failures.
@@
    Diagnosis. We empirically identify two coupled mechanisms behind MBPO’s collapses on DMC: (a) output-target scale mismatch (next-state vs. reward) that suppresses reward learning and induces critic underestimation and (b) variance inflation from residual targets that increases synthetic transition error and destabilizes MBPO’s policy improvements.
@@
More broadly, our results provide evidence that although improving average performance across larger benchmark sets is necessary for building trust, averages can hide critical flaws. After all, RL practitioners expect algorithms to work, with modest tuning, on specific problems, and average gains are of little value if we cannot explain or anticipate failures in particular tasks. TD3 (Fujimoto et al., 2018) provides a convenient example: it is known to perform well with dense rewards, fail in sparse settings, and succeed under sparse rewards when combined with Hindsight Experience Replay (Andrychowicz et al., 2018). This reflects an understanding of how structural properties of the underlying MDP (reward sparsity) shape algorithmic success. In much the same way, our findings underscore this need by providing a similar mapping: interactions between dynamics and reward scales, together with model training choices, can determine whether MBPO succeeds or fails even across seemingly similar tasks. Alongside higher average performance across diverse benchmarks, building up a body of prescriptive task–algorithm mappings is a necessary step toward ensuring RL methods are both effective and deployable in practice.
@@ 2.1 Reinforcement Learning, Model-Based RL, and Dyna-style Algorithms
On the OpenAI Gym benchmark suite, MBPO consistently demonstrates strong performance and high sample efficiency, often outperforming its model-free counterparts. This success has made MBPO the foundation for a wide range of subsequent work (Lai et al., 2020; 2022; Zheng et al., 2023; Wang et al., 2023; Li et al., 2024; Dong et al., 2024).
@@ 3.1 Generalization Issues In Dyna-style Reinforcement Learning
Benchmark-driven evaluations are the de facto standard in reinforcement learning, yet their validity as tests of generalization are rarely questioned. Can We Hop in General? (Voelcker et al., 2024) shows that benchmark choice can drastically alter conclusions, highlighting with Hopper environments that even different implementations of the same intuitive task yield contradictory results. Algorithms such as SAC (Haarnoja et al., 2019), MBPO (Janner et al., 2019), ALM (Ghugare et al., 2023), and DIAYN (Eysenbach et al., 2019) all report successes on some Hopper variants but fail on others, despite ostensibly targeting the same control problem. This reveals a broader issue: RL benchmarks are often treated as proxies for general capability, yet they may not even be representative of each other.
Building on this critique, Stealing That Free Lunch (STFL) (Barkley & Fridovich-Keil, 2025) systematically studied two Dyna-style methods, MBPO and ALM, across six OpenAI Gym and fifteen DeepMind Control Suite (DMC) tasks. Although all of these Gym and DMC tasks use MuJoCo (Todorov et al., 2012) and feature similar continuous control problems, MBPO and ALM’s results diverged sharply: they consistently outperformed their base algorithms in Gym, yet failed in seven of fifteen DMC environments, often failing to outperform even a random initial policy. In the case of MBPO, SAC not only outperformed MBPO across most DMC tasks, it did so with substantially less wall-clock time and compute. STFL investigated possible explanations — including overestimation bias, network plasticity, model fidelity, and extensive hyperparameter tuning — but found that none resolved the discrepancy. Crucially, it showed that introducing any synthetic rollout data into training consistently degraded policy performance.
Together, Can We Hop in General? (Voelcker et al., 2024) and Stealing That Free Lunch (Barkley & Fridovich-Keil, 2025) reveal how strongly RL results depend on benchmark choice and show that existing remedies fail to close the gap. Motivated by these findings, we turn to a focused analysis of MBPO to ask where, when, and why synthetic data causes its failures.
@@ 4 When, Where, and Why Synthetic Data Fails in MBPO
To enable this analysis, we build directly on STFL’s efficient reimplementation of MBPO, which replicated Gym results while simplifying rollouts to a one-step regime (Barkley & Fridovich-Keil, 2025). This implementation eliminates the potential confounder of accumulating model errors in longer-horizon rollouts, allowing us to attribute performance gaps in DMC to other factors.
@@ 4.3 Revisiting Residual Prediction: A Suboptimal Target for MBPO
This anomaly led us to revisit the standard practice of residual prediction, where models learn $s^{\prime}-s$ rather than $s^{\prime}$ directly. Residual prediction was popularized in early model-based RL approaches (Deisenroth & Rasmussen, 2011; Chua et al., 2018) and remains the default in modern world models (Hafner et al., 2021; 2023). The rationale is compelling: most state variables change slowly, so differences provide a strong inductive bias, allowing the model to focus on sparse dynamics rather than relearn the entire state at each step. This bias is particularly effective in smooth systems or for image-based states where consecutive frames are highly correlated (Sekar et al., 2020).
However, residual prediction is not universally superior. Its effectiveness depends strongly on system dynamics and noise characteristics (Lambert et al., 2018). In unstable regimes, residuals can amplify small errors that compound exponentially, producing rapid divergence. These results motivate the hypothesis that in DMC tasks, residual prediction inflates variance in the learned dynamics, yielding highly uncertain synthetic data that destabilizes MBPO’s training.
@@
Figure 3: (3(a)) Variance loss comparison between walker2d and humanoid-stand, showing substantially higher variance in the latter (1 seed). (3(b)) Predicted variance under residual versus direct next-state modeling, where direct prediction yields lower variance and a more stable fit (1 seed). (3(c)) Online returns for three ablation settings on humanoid-stand, demonstrating that only combining target normalization with direct prediction enables consistent policy improvement (3 seeds, mean ± std).
@@ 5.1 Ablations: Do we really need both?
To test whether the remedies identified in Sections 4.2 and 4.3 generalize beyond a single case, we performed ablations across the seven DMC tasks that MBPO previously failed in, as shown in Appendix E. FTFL succeeds in five of seven cases, enabling stable policy improvement where MBPO had previously failed. The two remaining environments, where both MBPO and STFL fail, are discussed in detail in Section 5.3. These results confirm that the failures identified earlier were not isolated artifacts but systemic issues that FTFL resolves in the majority of cases.
@@ 5.2 Generalization and Restoring the Promise of MBPO
Having established that FTFL resolves the coupled failures identified in Section 4, we evaluate it on the seven challenging DMC tasks where MBPO previously showed no improvement over a random policy. As shown in Figures 1 and 4, the gains are substantial. FTFL enables significant policy improvement in five of the seven tasks (humanoid-stand, humanoid-walk, humanoid-run, quadruped-walk, quadruped-run), consistently surpassing SAC and delivering reliable learning where MBPO collapsed. Furthermore, increasing the capacity of the model ensemble allows FTFL to surpass SAC by an even larger margin in the quadruped tasks, indicating the scalability of our approach.
@@
Figure 4: Return performance on six DMC tasks. Following Agarwal et al. (2022), solid lines show interquartile mean (IQM) returns aggregated across six seeds, and shaded regions denote 95% bootstrapped confidence intervals. Results are shown for FTFL with the original model and for Tuned FTFL with increased model capacity (effective only on humanoid tasks). One result was omitted for brevity, but raw averaged returns are provided for all seven tasks in Appendix D.
Figure 5: Return performance on six OpenAI Gym tasks, using the same IQM and confidence interval conventions as Figure 4. Raw averaged returns are provided in Appendix D.
@@ 5.3 Remaining Failures and Towards a Taxonomy of RL Failure Modes
While FTFL resolves the failure modes in quadruped and humanoid environments, it does not fully mitigate MBPO’s issues on hopper-hop and hopper-stand. In these Hopper tasks, scale and residual issues are corrected, reward collapse is prevented, and variance is reduced, consistent with the results in Section 5.1. Yet performance still stalls. An additional experiment (cf. Appendix F) shows that removing the contact sensor in the hopper observation (a discontinuous contact-force signal) allows FTFL to match SAC on hopper-stand, but not on hopper-hop, pointing to a modality-specific challenge unique to Hopper that was absent in other failing DMC environments.
@@ 6 Broader Implications and Conclusion
7 Acknowledgments
This work was supported by the National Science Foundation under Grant No. 2409535.
Appendix A Usage of Large Language Models (LLMs)
LLMs were used for paper writing assistance and to aid in brainstorming, software development, and experiment design.
Appendix B Reproducibility Statement
We have taken several steps to ensure the reproducibility of our results. Implementation details for SAC, MBPO, and FTFL are provided in Appendix C, with complete hyperparameter tables included in the appendix. Additional raw return curves, remediation ablations, and environment-specific analyses are reported in Appendix D, Appendix E, and Appendix F, respectively. Our results build on the efficient MBPO reimplementation from Barkley & Fridovich-Keil (2025), which is publicly available.
Appendix C SAC and MBPO Implementation Details
Our FTFL implementation builds directly on the codebase of Barkley & Fridovich-Keil (2025) (github.com/CLeARoboticsLab/STFL), using their default hyperparameters as reproduced below. In both FTFL and MBPO, input normalization is recomputed before every ensemble retraining by calculating independent means and standard deviations for the actions and states in the replay buffer. These statistics are stored and used to unit-normalize transitions sampled for training batches. Target normalization is performed in the same way, independently for rewards and next states at the same cadence, and is additionally required to denormalize model outputs back into the original next-state and reward scales.
@@ Appendix C SAC and MBPO Implementation Details
Appendix D Raw Return Curves for SAC, MBPO, FTFL, and Tuned FTFL in Gym and DMC
@@
Appendix E Remediation Ablation
@@
Appendix F FTFL and SAC Performance in Hopper Environments Without Touch Sensor Observations

### v3

# A Forensic Analysis of Synthetic Data in RL: Diagnosing and Solving Algorithmic Failures in Model-Based Policy Optimization

@@ 1 Introduction
The “no free lunch” principle highlights that no optimization method achieves universal optimality across every problem instance. In reinforcement learning (RL), this underscores that an algorithm’s effectiveness is shaped by the specific environment and its characteristics. For Dyna style model-based RL (MBRL) algorithms, wherein synthetic model-generated data is intended to improve sample efficiency and enable rapid policy improvement, recent evidence suggests that these benefits are not guaranteed. In Stealing That Free Lunch (STFL) (Barkley and Fridovich-Keil, 2025), Model Based Policy Optimization (MBPO) (Janner et al., 2019), a widely used and well cited Dyna style algorithm, was shown to match or exceed its model free counterpart, Soft Actor Critic (SAC) (Haarnoja et al., 2019), in OpenAI Gym (Brockman et al., 2016). However, in many DeepMind Control Suite (DMC) (Tassa et al., 2020) tasks, MBPO then failed to improve beyond a randomly initialized policy when trained from scratch, despite both benchmarks sharing the same MuJoCo physics engine (Todorov et al., 2012) and featuring similar high level task reward structures.
While STFL established that these failures exist, it did not explain why MBPO collapses in specific settings, other than noting the central role of synthetic data in driving failures. In this work, we close that gap. We focus on challenging DMC tasks where MBPO failed completely in STFL, often performing no better than a random initial policy throughout training (Barkley and Fridovich-Keil, 2025). Through targeted experiments, we identify two coupled mechanisms driving these failures.
@@
    Diagnosis. We empirically identify two coupled mechanisms behind MBPO’s collapses in DMC: (a) output-target scale mismatch (next-state vs. reward) that suppresses reward learning and induces critic underestimation and (b) variance inflation from residual targets that increases synthetic transition error and destabilizes MBPO’s policy improvements.
@@
More broadly, our results provide evidence that although improving average performance across larger benchmark sets is necessary for building trust, averages can hide critical flaws. After all, RL practitioners expect algorithms to work, with modest tuning, on specific problems, and average gains are of little value if we cannot explain or anticipate failures in particular tasks. TD3 (Fujimoto et al., 2018) provides a convenient example: it is known to perform well with dense rewards, fail in sparse settings, and succeed under sparse rewards when combined with Hindsight Experience Replay (Andrychowicz et al., 2017). This reflects an understanding of how structural properties of the underlying MDP (reward sparsity) shape algorithmic success. In much the same way, our findings underscore this need by providing a similar mapping: interactions between dynamics and reward scales, together with model training choices, can determine whether MBPO succeeds or fails even across seemingly similar tasks. Alongside higher average performance across diverse benchmarks, building up a body of prescriptive task–algorithm mappings is a necessary step toward ensuring RL methods are both effective and deployable in practice.
@@ 2.1 Reinforcement Learning, Model-Based RL, and Dyna-style Algorithms
On the OpenAI Gym benchmark suite, MBPO consistently demonstrates strong performance and high sample efficiency, often outperforming its model-free counterparts. This success has made MBPO the foundation for a wide range of subsequent work (Lai et al., 2020, 2021; Zheng et al., 2023; Wang et al., 2023; Li et al., 2024; Dong et al., 2024).
@@ 3.1 Generalization Issues In Dyna-style Reinforcement Learning
Benchmark-driven evaluations are the de facto standard in reinforcement learning, yet their validity as tests of generalization are rarely questioned. Can We Hop in General? (Voelcker et al., 2024) shows that benchmark choice can drastically alter conclusions, highlighting with Hopper environments that even different implementations of the same intuitive task yield contradictory results. Algorithms such as SAC (Haarnoja et al., 2019), MBPO (Janner et al., 2019), ALM (Ghugare et al., 2022), and DIAYN (Eysenbach et al., 2019) all report successes on some Hopper variants but fail on others, despite ostensibly targeting the same control problem. This reveals a broader issue: RL benchmarks are often treated as proxies for general capability, yet they may not even be representative of each other.
Building on this critique, Stealing That Free Lunch (STFL) (Barkley and Fridovich-Keil, 2025) systematically studied two Dyna-style methods, MBPO and ALM, across six OpenAI Gym and fifteen DeepMind Control Suite (DMC) tasks. Although all of these Gym and DMC tasks use MuJoCo (Todorov et al., 2012) and feature similar continuous control problems, MBPO and ALM’s results diverged sharply: they consistently outperformed their base algorithms in Gym, yet failed in seven of fifteen DMC environments, often failing to outperform even a random initial policy. In the case of MBPO, SAC not only outperformed MBPO across most DMC tasks, it did so with substantially less wall-clock time and compute. STFL investigated possible explanations — including overestimation bias, network plasticity, model fidelity, and extensive hyperparameter tuning — but found that none resolved the discrepancy. Crucially, it showed that introducing any synthetic rollout data into training consistently degraded policy performance.
Together, Can We Hop in General? (Voelcker et al., 2024) and Stealing That Free Lunch (Barkley and Fridovich-Keil, 2025) reveal how strongly RL results depend on benchmark choice and show that existing remedies fail to close the gap. Motivated by these findings, we turn to a focused analysis of MBPO to ask where, when, and why synthetic data causes its failures.
@@ 4 When, Where, and Why Synthetic Data Fails in MBPO
To enable this analysis, we build directly on STFL’s efficient reimplementation of MBPO, which replicated Gym results while simplifying rollouts to a one-step regime (Barkley and Fridovich-Keil, 2025). This implementation eliminates the potential confounder of accumulating model errors in longer-horizon rollouts, allowing us to attribute performance gaps in DMC to other factors.
@@ 4.2 Inappropriate Model Target Scale Leads to Severe Critic Underestimation
Figure 3: (3(a)) Variance loss comparison between walker2d and humanoid-stand, showing substantially higher variance in the latter (1 seed). (3(b)) Predicted variance under residual versus direct next-state modeling in the humanoid-stand task, where direct prediction yields lower variance and a more stable fit (1 seed). (3(c)) Online returns for three ablation settings on humanoid-stand, demonstrating that only combining target normalization with direct prediction enables consistent policy improvement (3 seeds, mean ± std).
@@ 4.3 Revisiting Residual Prediction: A Suboptimal Target for MBPO
This anomaly led us to revisit the standard practice of residual prediction, where models learn $s^{\prime}-s$ rather than $s^{\prime}$ directly. Residual prediction was popularized in early model-based RL approaches (Deisenroth and Rasmussen, 2011; Chua et al., 2018) and remains the default in modern world models (Hafner et al., 2020, 2023). The rationale is compelling: most state variables change slowly, so differences provide a strong inductive bias, allowing the model to focus on sparse dynamics rather than relearn the entire state at each step. This bias is particularly effective in smooth systems or for image-based states where consecutive frames are highly correlated (Sekar et al., 2020).
However, residual prediction is not universally superior. Its effectiveness depends strongly on system dynamics and noise characteristics (Lambert et al., 2022). In unstable regimes, residuals can amplify small errors that compound exponentially, producing rapid divergence. These results motivate the hypothesis that in DMC tasks, residual prediction inflates variance in the learned dynamics, yielding highly uncertain synthetic data that destabilizes MBPO’s training.
@@ 5.1 Ablations: Do we really need both?
To test whether the remedies identified in Sections 4.2 and 4.3 generalize beyond a single case, we performed ablations across the seven DMC tasks that MBPO previously failed in, as shown in Appendix C. FTFL succeeds in five of seven cases, enabling stable policy improvement where MBPO had previously failed. The two remaining environments, where both MBPO and STFL fail, are discussed in detail in Section 5.3. These results confirm that the failures identified earlier were not isolated artifacts but systemic issues that FTFL resolves in the majority of cases.
@@ 5.2 Generalization and Restoring the Promise of MBPO
Having established that FTFL resolves the coupled failures identified in Section 4, we evaluate it on the seven challenging DMC tasks where MBPO previously showed no improvement over a random policy. As shown in Figures 1 and 4, the gains are substantial. FTFL enables significant policy improvement in five of the seven tasks (humanoid-stand, humanoid-walk, humanoid-run, quadruped-walk, and quadruped-run), consistently surpassing SAC and delivering reliable learning where MBPO collapsed. Furthermore, increasing the capacity of the model ensemble allows FTFL to surpass SAC by an even larger margin in the quadruped tasks, indicating the scalability of our approach.
@@
Figure 4: Return performance on six DMC tasks. Following Agarwal et al. (2021), solid lines show interquartile mean (IQM) returns aggregated across six seeds, and shaded regions denote 95% bootstrapped confidence intervals. Results are shown for FTFL with the original model and for Tuned FTFL with increased model capacity (effective only on humanoid tasks). One result was omitted for brevity, but raw averaged returns are provided for all seven tasks in Appendix B.
Figure 5: Return performance on six OpenAI Gym tasks, using the same IQM and confidence interval conventions as Figure 4. Raw averaged returns are provided in Appendix B.
@@ 5.3 Remaining Failures and Towards a Taxonomy of RL Failure Modes
While FTFL resolves the failure modes in quadruped and humanoid environments, it does not fully mitigate MBPO’s issues on hopper-hop and hopper-stand. In these Hopper tasks, scale and residual issues are corrected, reward collapse is prevented, and variance is reduced, consistent with the results in Section 5.1. Yet performance still stalls. An additional experiment (cf. Appendix D) shows that removing the contact sensor in the hopper observation (a discontinuous contact-force signal) allows FTFL to match SAC on hopper-stand, but not on hopper-hop, pointing to a modality-specific challenge unique to Hopper that was absent in other failing DMC environments.
@@ 6 Broader Implications and Conclusion
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A SAC and MBPO Implementation Details
Our FTFL implementation builds directly on the codebase of Barkley and Fridovich-Keil (2025) (github.com/CLeARoboticsLab/STFL), using their default hyperparameters as reproduced below. In both FTFL and MBPO, input normalization is recomputed before every ensemble retraining by calculating independent means and standard deviations for the actions and states in the replay buffer. These statistics are stored and used to unit-normalize transitions sampled for training batches. Target normalization is performed in the same way, independently for rewards and next states at the same cadence, and is additionally required to denormalize model outputs back into the original next-state and reward scales.
@@ Appendix A SAC and MBPO Implementation Details
Appendix B Raw Return Curves for SAC, MBPO, FTFL, and Tuned FTFL in Gym and DMC
@@
Appendix C Remediation Ablation
@@
Appendix D FTFL and SAC Performance in Hopper Environments Without Touch Sensor Observations